THE ASYMPTOTIC NORMALITY OF AN INTERMEDIATE ORDER STATISTIC OF THE RANGES OF SUB-SAMPLES

J. W. H. SWANEPOEL

Department of Statistics and O.R., Potchefstroom University for CHE, Potchefstroom 2520, South Africa

(Received March 28, 1991; revised April 13, 1992)

Abstract. An intermediate order statistic based on the ranges of sub-samples is introduced and its limiting distribution derived under certain regularity conditions.

Key words and phrases: Intermediate order statistic, sample range, weak convergence.